# Cosmic dynamics and observational constraints in $f(Q)$ gravity with affine equation of state

### v1

@@ Abstract
In this study, we explore the cosmological implications of $f(Q)$ gravity with an affine equation of state (EoS). We study the impact of affine EoS on the cosmic evolution in model and derive the functional form of $f(Q)$ gravity realizing this kind of scenario. We show that the non-metricity scalar may describe the unified scenario in which the universe transits from the decelerated expansion into the accelerated expansion. The model parameters are constrained through the Bayesian analysis based on $\chi^{2}$ minimization technique with the observational data of the cosmic chronometer and supernovae type Ia. The cosmic dynamics based on constrained parameters describe the observable universe evolution. The present day values of the cosmological parameters along with the current age of the universe are compatible with the observations.
@@ 1 Introduction
The validation of accelerated cosmic expansion signifies the beginning of a new era in cosmological study. This phenomena was initially noticed by while observing type Ia supernovae [1, 2]. Subsequently, the universe’s expansion has been confirmed by numerous observations [3, 4, 5]. An unusual fluid/field with negative pressure is a cause of the universe’s late-time accelerated expansion [6, 7]. This kind of fluid/field has been termed as the ‘dark energy’ whose exact nature remains mysterious [8, 6]. In other words, the dark energy, theoretically modeled to generate negative pressure has been believed to be responsible for the universe’s accelerated expansion [9, 10, 11, 12, 13, 14]. The most widely accepted model of universe is the $\Lambda$ cold dark matter ($\Lambda$CDM) model. In this model, the dark energy is described by the constant vacuum energy density which is visualized as the cosmological constant in the model [6, 7]. This model is consistent with the observations, however, it faces several challenges [15, 6, 7, 13]. The enormous difference between the theoretically calculated and observational required values of vacuum energy corresponding to cosmological constant causes the issue of fine tuning [6, 7]. In order to resolve the challenges of $\Lambda$CDM model, the alternative approaches where the dark energy is a dynamical one is generally explored in the cosmological modeling.
@@
In the context of modified gravity theories, Jimenez et al. [33] proposed a gravity theory based on non-metricity and is termed as $f(Q)$ gravity. This theory is based on an approach which considers a variation of the symmetric teleparallel equivalence to General Relativity, here $Q$ is a non-metricity scalar. The studies [13, 34], suggest that the $f(Q)$ theory is one of the intriguing alternative gravity interpretations for explaining the cosmological evolution. Lazkoz et al. [35] investigated the accelerated expansion of the universe in context of its consistency with observational data. The exact cosmic solutions that are both anisotropic and isotropic were studied by Esposito et al.[36]. Harko et al. [37] developed a class of $f(Q)$ theories where $Q$ is non-minimally associated with the matter Lagrangian. In connection with the $f(Q)$ gravity theory, numerous further studies [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] suggest that the theory may have interesting cosmological and astrophysical implications, for review see [14].
@@
In this paper, we study the implications of $f(Q)$ gravity with the dark fluid satisfying affine equation of state. We search for an answer whether the $f(Q)$ gravity may have fluid description or not? In particular, for an affine equation of state of dark fluid, we derive the functional formulation of $f(Q)$ gravity. The issue of observational compatibility of model may describe the observational constraints on model parameters. The dark fluid have been studied in the isotropic and anisotropic spacetime with the General relativity framework [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71]. The affine equation of state describes the hydro-dynamically stable as well as unstable fluid propagation. The speed of sound remains constant for the fluid. The stable dark energy evolution may be visualized with the fluid satisfying affine equation of state given by $p_{d}=\alpha\rho_{d}-\rho_{0}$, where $\alpha$ and $\rho_{0}$ are constants. The classical stability criterion constrains $0\leq\alpha\leq 1$. In a non-conservative theory, this kind of EoS may naturally arise near the bouncing instant of bouncing universe [66, 68]. The low energy limits of this kind of EoS may have the $\Lambda$CDM limit [60]. Based on these peculiar properties, we aim to study the cosmological implications of this fluid in $f(Q)$ gravity to describes the fluid interpretation of this modified gravity.
@@
This paper is organized in the following way: An overview of the mathematical aspects of $f(Q)$ gravity in a flat FLRW background has been given in section (2). In Section (3), we obtain the cosmological solutions of the affine EoS model with $f(Q)$ gravity. In Section (4), the Bayesian statistical tools are used to investigate the compatibility of cosmological solution with the cosmic chronometer (CC) and supernovae type Ia (Pantheon) datasets. The cosmological dynamics based on the parameters such as the deceleration parameter, pressure, energy density, cosmographic parameter and the age of universe are studied in section (5). The conclusion with summary of results are given in Section (6).
@@ 2 An overview of $f(Q)$ gravity
Here, $f(Q)$ represents an arbitrary function of the nonmetricity scalar $Q$. The symbol $g$ denotes metric tensor’s determinant $g_{ab}$ and the Lagrangian density for matter is symbolized by $\mathcal{L}_{m}$. The quantity $Q$ can be characterized by two distinct traces as
@@
${}Q=-Q_{\sigma ab}P^{\sigma ab},$ (3)
Additionally, the superpotential tensor (which is a conjugate of non-metricity) can be introduced as:
${}4P^{\sigma}_{\phantom{\sigma}ab}=2Q_{(a\phantom{\sigma}b)}^{\phantom{\mu}\sigma}-\tilde{Q}^{\sigma}g_{ab}-Q^{\sigma}_{\phantom{\sigma}ab}-Q^{\sigma}g_{ab}-\delta^{\sigma}_{(a}Q_{b)}^{\phantom{\nu)}},$ (4)
@@
where, $f_{Q}=\frac{df(Q)}{dQ}$ and energy-momentum tensor may be written as $T_{\mu\nu}=-\frac{2}{\sqrt{-g}}\frac{\delta(\sqrt{-g}\mathcal{L}_{m})}{\delta g^{\mu\nu}}$. We are assuming the unit system where $18\pi G=c=$ 1.
@@
${}ds^{2}=a^{2}(t)\left(dx^{2}+dy^{2}+dz^{2}\right)-dt^{2},$ (6)
Under these conditions, the non-metricity scalar $Q$ may be described as $Q=6H^{2}$, where $H\equiv\frac{\dot{a}}{a}$ symbolizes the Hubble parameter and $a(t)$ is the scale factor. The overhead dot has been used to indicate the derivative with respect to cosmic time $t$. For the metric (6), the Friedmann equations can be expressed as follows [72]:
$H^{2}=\frac{1}{6f_{Q}}\left(\rho+\frac{f}{2}\right),\quad\dot{H}+3H^{2}+\frac{H\dot{f}_{Q}}{f_{Q}}=\frac{1}{2f_{Q}}\left(\frac{f}{2}-p\right)$ (7)
@@
$3H^{2}=\rho+\frac{F}{2}-QF_{Q},\quad\dot{H}(2QF_{QQ}+F_{Q}+1)+\frac{1}{4}\left(2QF_{Q}-F+Q\right)=-\frac{p}{2}$ (8)
The Eqs (8) may be seen as modifications of General Relativity equations, incorporating an additional term that arises from the non-metricity ($Q$) of space-time. This term may exhibit a characteristics similar to a fluid component associated with dark sector of the fluid component of universe, i.e. $p_{Q}=p_{d}$ and $\rho_{Q}=\rho_{d}$. Consequently, using Eqs. (8), we may write
$3H^{2}=(\rho+\rho_{d}),\quad 2\dot{H}+3H^{2}=-(p+p_{d}),$ (9)
@@
These equations may be used to obtain the solution of system and thus the resulting Hubble parameter may be used for the identification of the cosmic evolution in model. We proceed to setup the unified scenario of universe evolution in this $f(Q)$ model, where the transiting universe evolution may be described by the non-metricity scalar.
3 The cosmological solutions with the affine equation of state in $f(Q)$ model
We proceed by taking the fluid which may represent the hydro-dynamically dark energy. The fluid has been assumed to have form [56, 61, 71]
$p_{d}=\alpha\rho_{d}-\rho_{0}$ (11)
where $\alpha$ and $\rho_{0}$ are some constants. For $\rho_{0}=0$, one may have the usual relation of EoS as $p_{d}=\omega\rho_{d}$ for $\alpha=\omega$. The parameter $\alpha$ may be visualized as constant speed of sound and the classical stability criterion constrains $0\leq\alpha\leq 1$. The causality and stability issues are important as far as the physical acceptability of a cosmological model is concerned [73, 74]. In a given medium, the sound speed describes the propagation speed of the acoustic (pressure) waves. The behavior of squared sound speed may be described by ${c_{s}}^{2}=\frac{\partial{p_{d}}}{\partial{\rho_{d}}}$. The uncontrolled growth of energy density perturbations may be prevented by ${c_{s}}^{2}>0$. The prevention of causality violation would lead to ${c_{s}}^{2}<1$. In the subsequent sections, we consider the cases where $\rho=0$ and $\rho\neq 0$ in the Eqs. (9) to write the cosmological solutions. With these assumptions, we check the impact of $\rho$ term in cosmic dynamics yielded in these different models. The energy conservation equation for the fluid may be written as
$\dot{\rho}_{d}+3H(\rho_{d}+p_{d})=0$ (12)
One may observe that for $\rho_{d}<\frac{\rho_{0}}{1+\alpha}$, one may have the phantom evolution of dark energy in model. For $\rho_{d}=\frac{\rho_{0}}{1+\alpha}$, the dynamics of constant vacuum energy may be visualized. For $\rho_{d}>\frac{\rho_{0}}{1+\alpha}$, the null energy condition will be satisfied in model, as a consequence the fluid distribution may either have the quintessence domination or the normal matter (composed of matter or radiation) domination. Based on these possible evolution dynamics, we probe the universe evolution under influence of dark fluid (given by Eq. (11)) in $f(Q)$ model.
@@ 3.1 Model I
For the case $\rho=0$ in Eqs. (9) and (10), we may obtain $\rho_{d}=3H^{2}$ and it would lead to $F(Q)=-Q+C\sqrt{Q}$, where $C$ is an integration constant. The form of $f(Q)$ function may be written as
$f(Q)=C\sqrt{Q}$ (13)
The above functional form of $f(Q)$ may describe the evolution with dark fluid having form (11). The continuity equation (12) may lead to the form of $\rho_{d}$ depending on the scale factor as
${}\rho_{d}=\frac{\rho_{0}+(1+z)^{3(\alpha+1)}}{\alpha+1}$ (14)
This Eq. (14) would yield the Hubble parameter for this case as
${}H(z)=H_{0}\sqrt{\frac{(1+z)^{3(1+\alpha)}+\rho_{0}}{\rho_{0}+1}},\quad\text{where,}\quad H_{0}=\sqrt{\frac{\rho_{0}+1}{3(1+\alpha)}}$ (15)
The above Hubble parameter may describe the cosmic evolution in the $f(Q)$ gravity yielded by the fluid following affine equation of state (11). The observational constraints on the model parameters $\alpha,\rho_{0}$ and $H_{0}$ may describe the observational compatibility of model along with the validity of causality principle in this model.
@@ 3.2 Model II
In the present case, we take $\rho\neq 0$ in equation (9) and assume that $\rho$ is the energy density of pressure-less matter satisfying $p=0$. As a consequence, $\rho=\rho_{m0}a^{-3}$, where $\rho_{m0}$ is energy density of matter at $z=0$. In this case, $\rho_{d}$ may be the energy density of dark energy. In this case, the Hubble parameter may be obtained as
${}H(z)=H_{0}\sqrt{(1+z)^{3}\left[\Omega_{m0}+\left(\frac{1-\Omega_{m0}}{\rho_{0}+1}\right)(1+z)^{3\alpha}\right]+\left(\frac{1-\Omega_{m0}}{\rho_{0}+1}\right)\rho_{0}}$ (16)
where $\Omega_{m0}=\frac{\rho_{m0}}{3H_{0}^{2}}$. This Hubble parameter may describe the cosmic dynamics in $f(Q)$ gravity where the fluids are composed of the pressure-less matter and dark energy satisfying $p=0$ and $p_{d}=\alpha\rho_{d}-\rho_{0}$ respectively.

### v2

@@ Abstract
In this study, we explore the cosmological implications of $f(Q)$ gravity with the barotropic fluid satisfying the affine equation of state (EoS). We study the impact of affine EoS on the cosmic evolution in the model and derive the functional form $f(Q)$ in this gravity. We show that the non-metricity scalar may describe the unified scenario in which the universe transits from the decelerated expansion into the accelerated expansion. The model parameters are constrained by the Bayesian analysis based on $\chi^{2}$ minimization technique with the observational data of the cosmic chronometer and supernovae type Ia. The cosmic dynamics based on constrained parameters describe the observable universe evolution. The present day values of the cosmological parameters along with the current age of the universe are compatible with the observations.
@@ 1 Introduction
The validation of accelerated cosmic expansion signifies the beginning of a new era in cosmological study. This phenomena was initially noticed while observing the type Ia supernovae [1, 2]. Subsequently, the universe’s expansion has been confirmed by numerous observations [3, 4, 5]. An unusual fluid/field with negative pressure is a cause of the universe’s late-time accelerated expansion [6, 7]. This kind of fluid/field has been termed as the ‘dark energy’ whose exact nature remains mysterious [8, 6]. In other words, the dark energy, theoretically modeled to generate negative pressure has been believed to be responsible for the universe’s accelerated expansion [9, 10, 11, 12, 13, 14]. The most widely accepted model of the universe is the $\Lambda$ cold dark matter ($\Lambda$CDM) model. In this model, the dark energy is described by the constant vacuum energy density which is visualized as the cosmological constant in the model [6, 7]. This model is consistent with the observations, however, it faces several challenges [15, 6, 7, 13]. The enormous difference between the theoretical and observational values of vacuum energy corresponding to cosmological constant causes the issue of fine tuning [6, 7]. In order to resolve the challenges of the $\Lambda$CDM model, the alternative approaches where the dark energy is a dynamical one is generally explored in cosmological modeling.
@@
In the context of modified gravity theories, Jimenez et al. [33] proposed a gravity theory based on non-metricity and is termed as $f(Q)$ gravity. This theory is based on an approach which considers a variation of the symmetric teleparallel equivalence to General Relativity, here $Q$ is a non-metricity scalar. The studies [13, 34], suggest that the $f(Q)$ theory is one of the intriguing alternative gravity interpretations for explaining cosmological evolution. Lazkoz et al. [35] investigated the accelerated expansion of the universe in context of its consistency with observational data. The exact cosmic solutions that are both anisotropic and isotropic were studied by Esposito et al. [36]. Harko et al. [37] developed a class of $f(Q)$ theories where $Q$ is non-minimally associated with the matter Lagrangian. In connection with the $f(Q)$ gravity theory, numerous further studies [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59] suggest that the theory may have interesting cosmological and astrophysical implications, for review see [14].
@@
In this paper, we study the implications of $f(Q)$ gravity with the dark fluid satisfying affine equation of state. We search for an answer whether the $f(Q)$ gravity may have fluid description or not? In particular, for an affine equation of state of dark fluid, we derive the functional formulation of $f(Q)$ gravity. The issue of observational compatibility of model may imply for the observational constraints on model parameters. The dark fluid have been studied in the isotropic and anisotropic spacetime with the General relativity framework [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]. The affine equation of state describes the hydro-dynamically stable as well as unstable fluid propagation. The speed of sound remains constant for the fluid. The stable dark energy evolution may be visualized with the fluid satisfying affine equation of state given by $p_{d}=\alpha\rho_{d}-\rho_{0}$, where $\alpha$ and $\rho_{0}$ are constants. The classical stability criterion constrains $0\leq\alpha\leq 1$. In a non-conservative theory, this kind of EoS may naturally arise near the bouncing instant of the bouncing universe [70, 72]. The low energy limits of this kind of EoS may have the $\Lambda$CDM limit [64]. Based on these peculiar properties, we aim to study the cosmological implications of this dark fluid in $f(Q)$ gravity to describe the fluid interpretation of this gravity.
@@
This paper is organized in the following way: An overview of the mathematical aspects of $f(Q)$ gravity in a flat FLRW background has been given in section (2). In Section (3), we obtain the cosmological solutions of the affine EoS model with $f(Q)$ gravity. In Section (4), the Bayesian statistical tools are used to investigate the compatibility of cosmological solutions with the cosmic chronometer (CC) and supernovae type Ia (Pantheon) datasets. The cosmological dynamics based on the parameters such as the deceleration parameter, pressure, energy density, cosmographic parameter and the age of the universe are studied in section (5). The conclusion with summary of results are given in Section (6).
@@ 2 An overview of $f(Q)$ gravity
Here, $f(Q)$ represents an arbitrary function of the non-metricity scalar $Q$. The symbol $g$ denotes metric tensor’s determinant $g_{ab}$ and the Lagrangian density for matter is symbolized by $\mathcal{L}_{m}$. The quantity $Q$ can be characterized by two distinct traces as
@@
${}Q=-Q_{\sigma ab}P^{\sigma ab}.$ (3)
Additionally, the super-potential tensor (which is a conjugate of non-metricity) can be introduced as:
${}4P^{\sigma}_{\phantom{\sigma}ab}=2Q_{(a\phantom{\sigma}b)}^{\phantom{\mu}\sigma}-\tilde{Q}^{\sigma}g_{ab}-Q^{\sigma}_{\phantom{\sigma}ab}-Q^{\sigma}g_{ab}-\delta^{\sigma}_{(a}Q_{b)}^{\phantom{\nu)}}.$ (4)
@@
where, $f_{Q}\equiv\frac{df(Q)}{dQ}$ and, the energy-momentum tensor may be written as $T_{ab}=-\frac{2}{\sqrt{-g}}\frac{\delta(\sqrt{-g}\mathcal{L}_{m})}{\delta g^{ab}}$. We are assuming the unit system where $18\pi G=c=$ 1.
@@
${}ds^{2}=a^{2}(t)\left(dx^{2}+dy^{2}+dz^{2}\right)-dt^{2}.$ (6)
Under these conditions, the non-metricity scalar may be described as $Q=6H^{2}$ [33], where $H\equiv\frac{\dot{a}}{a}$ symbolizes the Hubble parameter and $a(t)$ is the scale factor. The overhead dot has been used to indicate the derivative with respect to cosmic time $t$. For the metric (6), the Friedmann equations can be expressed as [76]:
$H^{2}=\frac{1}{6f_{Q}}\left(\rho+\frac{f}{2}\right),\quad\dot{H}+3H^{2}+\frac{H\dot{f}_{Q}}{f_{Q}}=\frac{1}{2f_{Q}}\left(\frac{f}{2}-p\right).$ (7)
@@
$3H^{2}=\rho+\frac{F}{2}-QF_{Q},\quad\dot{H}(2QF_{QQ}+F_{Q}+1)+\frac{1}{4}\left(2QF_{Q}-F+Q\right)=-\frac{p}{2}.$ (8)
The Eq. (8) may be seen as modifications of the General Relativity equations, incorporating an additional term that arises from the non-metricity ($Q$) of space-time. This term may exhibit a characteristics similar to a fluid component associated with dark sector of the fluid component of universe, i.e. $p_{Q}=p_{d}$ and $\rho_{Q}=\rho_{d}$. Consequently, using Eqs. (8), we may write
$3H^{2}=(\rho+\rho_{d}),\quad 2\dot{H}+3H^{2}=-(p+p_{d}).$ (9)
@@
These equations may be used to obtain the solution of the system and thus the resulting Hubble parameter may be used for the identification of the cosmic evolution in the model. We proceed to set up the unified scenario of universe evolution in this $f(Q)$ model, where the transiting universe evolution may be described by the non-metricity scalar.
3 The cosmological solutions with the dark fluid in $f(Q)$ model
We proceed by taking the fluid which may represent the hydro-dynamically stable dark energy. The fluid has been assumed to have form [60, 65, 75]
$p_{d}=\alpha\rho_{d}-\rho_{0},$ (11)
where $\alpha$ and $\rho_{0}$ are some constants. For $\rho_{0}=0$, one may have the usual EoS $p_{d}=\omega\rho_{d}$ for $\alpha=\omega$. The parameter $\alpha$ may be visualized as constant speed of sound and the classical stability criterion constrains $\alpha$ as $0\leq\alpha\leq 1$. The causality and stability issues are important as far as the physical acceptability of a cosmological model is concerned [77, 78]. In a given medium, the sound speed describes the propagation speed of the acoustic (pressure) waves. The behavior of squared sound speed may be described by ${c_{s}}^{2}=\frac{\partial{p_{d}}}{\partial{\rho_{d}}}$. The uncontrolled growth of energy density perturbations may be prevented by ${c_{s}}^{2}>0$. The prevention of causality violation would lead to ${c_{s}}^{2}<1$.
In the subsequent sections, we study two different cases arising from Eq. (9). In general, the dark fluid may give rise to the transiting universe evolution (from matter dominated era into dark energy dominated era). We probe whether the dark fluid satisfying affine EoS yields this kind of scenario or not? For the affirmative case, we search for the functional form of $f(Q)$ in the $f(Q)$ gravity. In order to probe these issues, we assume $\rho=0$ in Eq. (9). As a consequence, the evolution of the Hubble parameter will be governed by the dark fluid (11) only. We nomenclate this considered scenario as Model I. In another case, we assume that the dark fluid (with energy density $\rho_{d}$ and pressure $p_{d}$) will account for the dark energy only and then assume that $\rho$ appears in the Eq. (9) is the matter energy density, that is $\rho=\rho_{m}$. As a consequence, $p=p_{m}=0$. These assumptions are motivated by the fact that matter (cold dark matter and baryonic matter) and dark energy are the dominant components of the present day universe. In the Model II, we probe for the universe’s evolution in $f(Q)$ gravity when the matter (following $p_{m}=0$) and dark energy (following Eq. (11)) are the dominant components of the universe in the model. We further describe the corresponding differences in universe evolution for these two cases. In other words, these two cases enable us to check the impact of the $\rho_{m}$ term in the cosmic dynamics yielded in these different models. The energy conservation equation for the dark fluid may be written as
$\dot{\rho}_{d}+3H(\rho_{d}+p_{d})=0.$ (12)
One may observe that for $\rho_{d}<\frac{\rho_{0}}{1+\alpha}$, one may have the phantom evolution of dark energy in model. For $\rho_{d}=\frac{\rho_{0}}{1+\alpha}$, the dynamics of constant vacuum energy may be visualized. For $\rho_{d}>\frac{\rho_{0}}{1+\alpha}$, the null energy condition will be satisfied in model, as a consequence the fluid distribution may either have the quintessence domination or the normal matter (composed of matter or radiation) domination. Based on these possible evolutionary dynamics, we probe the universe’s evolution under influence of dark fluid (given by Eq. (11)) in $f(Q)$ model.
@@ 3.1 Model I
In this model, we assume that the dark fluid (Eq. (11)) encompassed within $f(Q)$ gravity is the dominant component of the universe. For this case, $\rho=0$ in Eq. (9) leading to $\rho_{d}=3H^{2}$. From Eq. (11), it would lead to $F(Q)=-Q+C\sqrt{Q}$, where $C$ is an integration constant. The form of $f(Q)$ function may be written as
$f(Q)=C\sqrt{Q}.$ (13)
The above functional form of $f(Q)$ may describe the universe evolution yielded by the dark fluid (11). In other words, the affine EoS cosmology may belong to the class of cosmologies yielded by $f(Q)=C\sqrt{Q}$ gravity. The dark fluid conservation equation (12) may lead to the $\rho_{d}$ form depending on the scale factor as
${}\rho_{d}=\frac{\rho_{0}+(1+z)^{3(\alpha+1)}}{\alpha+1}.$ (14)
In this case, the Eqs. (9) and (14) would yield the Hubble parameter as
${}H(z)=H_{0}\sqrt{\frac{(1+z)^{3(1+\alpha)}+\rho_{0}}{\rho_{0}+1}},\quad\text{where,}\quad H_{0}=\sqrt{\frac{\rho_{0}+1}{3(1+\alpha)}}.$ (15)
The above Hubble parameter may describe the cosmic evolution in the $f(Q)$ gravity yielded by the fluid following affine equation of state (11). The observational constraints on the model parameters $\alpha,\rho_{0}$ and $H_{0}$ may describe the observational compatibility of the model along with the validity of causality principle in this model.
@@ 3.2 Model II
In this case, we take the universe composed of matter (cold dark matter and baryonic matter) with the dark energy governed by the dark fluid (11) in the $f(Q)$ gravity. For the present case, we take $\rho=\rho_{m}$ in Eq. (9), where $\rho_{m}$ denotes the energy density of matter satisfying EoS $p_{m}=0$. The conservation equation of matter $\dot{\rho}_{m}+3H\rho_{m}=0$ would lead to $\rho_{m}=\rho_{m0}a^{-3}$, where $\rho_{m0}$ is energy density of matter at $z=0$. And, $\rho_{d}$ is the dark energy density following Eq. (12). In this case, the Hubble parameter may be obtained as
${}H(z)=H_{0}\sqrt{\Omega_{m0}(1+z)^{3}+\Omega_{\Lambda 0}+\Omega_{a0}(1+z)^{3(1+\alpha)}}$ (16)
where $\Omega_{m0}\equiv\frac{\rho_{m0}}{3H_{0}^{2}}$, $\Omega_{\Lambda 0}\equiv\frac{\rho_{0}}{3(1+\alpha)H_{0}^{2}}$ and $\Omega_{a0}\equiv\frac{\rho_{\alpha 0}}{3(1+\alpha)H_{0}^{2}}$. In the above expression, we use $\frac{a_{0}}{a}=1+z$ with $a_{0}=1$ for present day universe. At $z=0$ in present era, $H=H_{0}$ with constraint $\Omega_{m0}+\Omega_{\Lambda 0}+\Omega_{a0}=1$. This Hubble parameter may describe the cosmic dynamics in $f(Q)$ gravity where the fluids are composed of matter and dark energy satisfying $p_{m}=0$ and $p_{d}=\alpha\rho_{d}-\rho_{0}$ respectively. In this study, we ignore the contribution of radiation since it is observationally very small in the late-time universe [5]. We probe the observational compatibility of these models with the observational data and derive the constraint on model parameters.